OpenStudy (gm100299):

Check my work PLEASE There are 9 red checkers and 3 black checkers in a bag. Checkers are selected one at a time, with replacement. Each time, the color of the checker is recorded. Find the probability of selecting a red checker exactly 8 times in 10 selections. Show your work. Red checker: 9/12 = 3/4 Black checker: 9/12 = 1/4 (3/4)(3/4)(3/4)(3/4)(3/4)(3/4)(3/4)(3/4)(1/4)(1/4) = (3/4)^8 (1/4)^2 10C8 = 10!/ (8!2!) = 10(9/2) = 45 P(8) = 45 (3/4)^8 (1/4)^2 = 0.281568 = 28%

Directrix (directrix):

Binomial Probability Problem with n = 10 and p(red) = 12/15 P(x = 8) = C( 10, 8) * (3/4)^8 * (1/4) ^ 3

Directrix (directrix):

Let me crank that out and compare with your answer.

Directrix (directrix):

My equation is messed up. Sorry.

Directrix (directrix):

Binomial Probability Problem with n = 10 and p(red) = 4/5 P(x = 8) = C( 10, 8) * (3/4)^8 * (1/4) ^2

Directrix (directrix):

Revised Equation ^^^^^

Directrix (directrix):

Probability = .28156 Is that what you got?

OpenStudy (gm100299):

Yes that's what I got too, thanks

Directrix (directrix):

.28156 Sorry, I left out the 1. http://www.wolframalpha.com/input/?i=C(+10,+8)+*+(3%2F4)%5E8+*+(1%2F4)+%5E2

Directrix (directrix):

Alrighty, then.